# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 9 Densification — Phase 2
Let $G$ be a graph on $n$ vertices. Let $X,Y,Z\subset V(G)$ be disjoint sets such that $|X|=|Y|=|Z|$ and $(X,Y)$ and $(Y,Z)$ are $\delta$-full, but $(Y,Z)$ is $\delta$-empty. Then $\operatorname{surp}(G)\geq(1/4-3\delta)|X|$.
@@
$\operatorname{surp}(G)\geq\operatorname{surp}(G^{\prime})=\mbox{mc}(G^{\prime})-\frac{e(G^{\prime})}{2}\geq|X|^{2}(2-2\delta)-\left(\frac{7}{4}+\frac{\delta}{2}\right)|X|^{2}=|X|^{2}\left(\frac{1}{4}-\frac{5}{2}\delta\right).$
@@
The first part follows from Alon and Shapira [7], so we only prove the second part. Let $\tilde{G}$ be the disjoint union of cliques that is $\delta$-close to $G$. Then each cherry of $G$ contains at least one edge or non-edge from $\tilde{G}\Delta G$, so we are done. ∎
@@
Let $s\in(0,1/4)$ and $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph such that $|\lambda_{n}|\leq n^{s}$. Then $G$ is $\delta$-close to the union of cliques.
@@
Let $\delta_{0}>0$ be specified later, depending only on $\delta$. Let $\varepsilon$ be the constant guaranteed by Lemma 9.1 with respect to $\delta_{0}$. We have
$\sum_{\lambda_{i}>0}\lambda_{i}=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq|\lambda_{n}|n\leq n^{1+s}.$
Also, by Lemma 6.3, we have $4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}$ for every $T\geq n^{1/2+s}\geq 2|\lambda_{n}|\sqrt{n}$. Hence, we can apply Lemma 6.7 to the sequence of positive eigenvalues with $q:=1/4>s$ to conclude that for every $H\leq n$, we have
$\sum_{0<\lambda_{i}<H}\lambda_{i}^{2}\leq O(n^{1+4s}H^{1-4s}).$
$\sum_{\lambda_{i}<H}\lambda_{i}^{2}\leq O_{s}(n^{1+4s}H^{1-4s})+n^{3/2}.$
Set $H=\varepsilon_{0}n$, where $\varepsilon_{0}>0$ is specified later. Then $n^{1+4s}H^{1-4s}=n^{2}\varepsilon_{0}^{1-4s}$. Hence, we can choose $\varepsilon_{0}$ (depending only on $s$ and $\varepsilon$) such that
$\sum_{\lambda_{i}<H}\lambda_{i}^{2}\leq\varepsilon n.$
Let $A$ be the adjacency matrix of $G$, then $A$ has $pn(n-1)>pn^{2}/2$ one entries, and we note that $p/2\geq 4\delta_{0}$ holds. Let
$B=\sum_{\lambda_{i}\geq H}\lambda_{i}v_{i}v_{i}^{T},$
then
$A-B||_{F}^{2}=\sum_{\lambda_{i}<H}\lambda_{i}^{2}\leq\varepsilon n^{2}.$
Furthermore, the rank of $B$ is $r:=N_{H}\leq\frac{n^{2}}{H^{2}}\leq\varepsilon_{0}^{-2}$, where in the first inequality we used that $n^{2}\geq\sum_{\lambda_{i}\geq H}\lambda_{i}^{2}\geq H^{2}N_{H}$. Hence, we can apply Lemma 9.1 to conclude that there is a $\delta_{0}$-regular partition $V_{1},\dots,V_{K}$ for some $K$ with $1/\delta_{0}<K<O_{r,\delta_{0}}(1)=O_{s,\delta}(1)$.
@@
By Lemma 9.2, we cannot have that $(V_{i},V_{j})$ and $(V_{j},V_{k})$ are $\delta_{0}$-full, but $(V_{i},V_{k})$ is $\delta_{0}$-empty. Therefore, each cherry belongs to one of the four categories. We observe that the number of cherries belonging to each category is at most $O(\delta_{0}n^{3})$. Indeed, for (i), there are $O(K^{2})$ choices for the set $\{i,j,k\}$, and then there are at most $(n/K)^{3}$ choices for $x,y,z$, so in total $O(K^{2}(n/K)^{3})=O(n^{3}/K)=O(\delta_{0}n^{3})$. For (ii), we use the fact that there are at most $\delta_{0}K^{2}$ non-$\delta_{0}$-homogeneous pairs $(V_{i},V_{j})$, so the number of choices for $(V_{i},V_{j},V_{k})$ is at most $O(\delta_{0}K^{3})$. For (iii) and (iv), we observe that if we fixed $(V_{i},V_{j},V_{k})$, then there are at most $\delta_{0}(n/K)^{3}$ choices for $x,y,z$. Indeed, if say $(V_{i},V_{j})$ is $\delta_{0}$-empty, the pair $(x,y)$ can be chosen from only the $\delta_{0}(n/K)^{2}$ edges between $V_{i}$ and $V_{j}$.
In conclusion, the number of cherries in $G$ is $O(\delta_{0}n^{3})$. But then by Lemma 9.3, $G$ is $O(\delta_{0})^{c}$-close to a disjoint union of cliques for some absolute constant $c>0$. We are done by setting $\delta_{0}>0$ sufficiently small with respect to $\delta$.
@@
Let $s\in(0,1/4)$, $p>0$ and $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density $p$ such that $|\lambda_{n}|\leq n^{s}$. Then $G$ contains a subgraph on at least $pn/2$ vertices of edge density at least $1-\delta$.
@@
Let $\delta_{0}=\delta p^{2}/2$. By Theorem 9.4, $G$ is $\delta_{0}$-close to some graph $H$ that is the disjoint union of cliques. Let $C_{1},\dots,C_{k}$ be the vertex sets of the cliques forming $H$, then
$e(H)=\sum_{i=1}^{k}\binom{|C_{i}|}{2}\leq\sum_{i=1}^{k}|C_{i}|^{2}\leq n\max_{i\in[k]}|C_{i}|.$
As $e(H)\geq e(G)-\delta_{0}n^{2}\geq pn^{2}/2$, this shows that at least one of the $C_{i}$’s has size at least $pn/2$. Without loss of generality, $|C_{1}|\geq pn/2$. But then $G[C_{1}]$ has at least $\binom{|C_{1}|}{2}-\delta_{0}n^{2}$ edges, so $G[C_{1}]$ has edge density at least $1-2\delta_{0}/p^{2}=1-\delta$. ∎
@@
Let $s\in(0,1/30)$, $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Then $G$ is $\delta$-close to a disjoint union of cliques.
@@
One of the key differences compared to the proof of Theorem 9.4 is that we use Lemma 6.5 to have $4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}$ satisfied for every $T\geq n^{1-\frac{1}{24}+\frac{s}{4}}$. Then setting $q=1/30$, the condition $s\in(0,1/30)$ ensures that $s<q$ and $1-q\geq 1-\frac{1}{24}+\frac{s}{4}$, so we can apply Lemma 6.7. Another difference is that we bound $\sum_{0<\lambda_{i}}\lambda_{i}^{2}$ using Lemma 5.3, (ii), which gives $\sum_{0<\lambda_{i}}\lambda_{i}^{2}\leq O(\sqrt{n}\operatorname{surp}^{*}(G))\leq O(n^{3/2+s})=o(n^{2})$. ∎
@@
Let $s\in(0,1/30)$, $p>0$ and $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density $p$ such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Then $G$ contains a subgraph on at least $pn/2$ vertices of edge density at least $1-\delta$.
@@ 10 Densification — Phase 3
In this section, we prove that graphs of edge density $1-\varepsilon$ and large smallest eigenvalue (or small surplus) must contain a subgraph of edge density $1-n^{-\varepsilon^{\prime}}$ on almost the same number of vertices. The proof of this proceeds via a density increment argument, as summarized in the following lemma.
@@
Let $s\in(0,1/3)$, and let $n$ be sufficiently large with respect to $s$. Let $G$ be an $n$-vertex graph with edge density 1p1-p, where $p<10^{-5}$. Assume that $|\lambda_{n}|<n^{s}$. Then $G$ contains an $n/2$-vertex induced subgraph of edge density at least $1-10^{7}p^{3}-O(n^{3s-1})$.
@@
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and corresponding orthonormal basis of eigenvectors $v_{1},\dots,v_{n}$. Define $B=A-\lambda_{1}v_{1}v_{1}^{T}$. The key idea of the proof is to consider the following triple Hadamard product:
$D=(B+|\lambda_{n}|I)^{\circ 3}=B^{\circ 3}+3|\lambda_{n}|B\circ B\circ I+3|\lambda_{n}|^{2}B\circ I+|\lambda_{n}|^{3}I.$
As $B+|\lambda_{n}|I=|\lambda_{n}|v_{1}v_{1}^{T}+\sum_{i=2}^{n}(\lambda_{i}+|\lambda_{n}|)v_{i}v_{i}^{T}$, we have that $B+|\lambda_{n}|I$ is positive semidefinite. Therefore, $D$ is also positive semidefinite by the Schur product theorem. Then, using the principal eigenvector $v_{1}$, we identify a set of well-behaved vertices $U$, and evaluate the product $0\leq\mathds{1}_{U}^{T}D\mathds{1}_{U}$.
We now give the details. Let $U$ be the set of vertices $i\in[n]$ that satisfy $v_{1}(i)\geq(1-8p)/\sqrt{n}$.
@@
$|U|\geq n/2$.

### v2

@@ 8 Densification — Phase 2
Let $G$ be a graph on $n$ vertices. Let $X,Y,Z\subset V(G)$ be disjoint sets such that $|X|=|Y|=|Z|$ and $(X,Y)$ and $(Y,Z)$ are $\delta$-full, but $(X,Z)$ is $\delta$-empty. Then $\operatorname{surp}(G)\geq(1/4-3\delta)|X|^{2}$.
@@
$\operatorname{surp}(G)\geq\operatorname{surp}(G^{\prime})\geq e\big(G[Y,X\cup Z]\big)-\frac{e(G^{\prime})}{2}\geq|X|^{2}(2-2\delta)-\left(\frac{7}{4}+\frac{\delta}{2}\right)|X|^{2}\geq|X|^{2}\left(\frac{1}{4}-3\delta\right).\hfill\qed$
@@
The first part follows from Alon and Shapira [8], so we only prove the second part. Let $\widetilde{G}$ be the disjoint union of cliques that is $\delta$-close to $G$. Then each cherry of $G$ contains at least one edge or non-edge from $\widetilde{G}\Delta G$, so we are done. ∎
@@
Let $\gamma\in(0,1/4)$, $\delta>0$, and let $n$ be sufficiently large. If $G$ is an $n$-vertex graph with $|\lambda_{n}|\leq n^{\gamma}$, then $G$ is $\delta$-close to the vertex-disjoint union of cliques.
@@
Let $\delta_{0}>0$ be a sufficiently small constant, depending only on $\delta$. Let $\varepsilon$ be the constant required so that Lemma 8.2 would hold with the parameter $\delta_{0}$. By Lemma 8.1, there is a matrix $B$ of rank $r=O_{\gamma,\varepsilon}(1)$ such that $\|A-B\|_{F}^{2}\leq\varepsilon n^{2}$. Hence, we can apply Lemma 8.2 to conclude that there is a $\delta_{0}$-regular partition $V_{1},\dots,V_{K}$ for some $K$ with $1/\delta_{0}<K<O_{r,\delta_{0}}(1)=O_{\gamma,\delta}(1)$.
@@
By Lemma˜8.3, we cannot have that $(V_{i},V_{j})$ and $(V_{j},V_{k})$ are $\delta_{0}$-full, but $(V_{i},V_{k})$ is $\delta_{0}$-empty. Therefore, each cherry belongs to one of the four categories. We observe that the number of cherries belonging to each category is at most $O(\delta_{0}n^{3})$. Indeed, for (i), there are $O(K^{2})$ choices for the set $\{i,j,k\}$, and then there are at most $(n/K)^{3}$ choices for $x,y,z$, so in total $O(K^{2}(n/K)^{3})=O(n^{3}/K)=O(\delta_{0}n^{3})$. For (ii), we use the fact that at most $\delta_{0}K^{2}$ pairs $(V_{i},V_{j})$ are not $\delta_{0}$-homogeneous to derive that the number of choices for $(V_{i},V_{j},V_{k})$ is $O(\delta_{0}K^{3})$. So the number of cherries belonging to (ii) is $O(\delta_{0}K^{3}(n/K)^{3})=O(\delta_{0}n^{3})$. For (iii) and (iv), we observe that if we fixed $(V_{i},V_{j},V_{k})$, then there are at most $\delta_{0}(n/K)^{3}$ choices for $x,y,z$. Indeed, if say $(V_{i},V_{j})$ is $\delta_{0}$-empty, the pair $(x,y)$ can be chosen from only the $\delta_{0}(n/K)^{2}$ edges between $V_{i}$ and $V_{j}$. So the number of cherries belonging to (iii) or (iv) is $O(\delta_{0}n^{3})$.
In conclusion, the number of cherries in $G$ is $O(\delta_{0}n^{3})$. But then by Lemma 8.4, $G$ is $O(\delta_{0})^{c}$-close to a disjoint union of cliques for some absolute constant $c>0$. We are done by setting $\delta_{0}>0$ sufficiently small with respect to $\delta$. ∎
@@
Let $\gamma\in(0,1/4)$, $p>0$ and $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density $p$ such that $|\lambda_{n}|\leq n^{\gamma}$. Then $G$ contains a subgraph on at least $pn/2$ vertices of edge density at least $1-\delta$.
@@
Let $\delta_{0}=\delta p^{2}/16$. By Theorem˜1.4, $G$ is $\delta_{0}$-close to some graph $H$ that is the disjoint union of cliques. Let $C_{1},\dots,C_{k}$ be the vertex sets of the cliques forming $H$, then
$e(H)=\sum_{i=1}^{k}\binom{|C_{i}|}{2}\leq\sum_{i=1}^{k}\frac{|C_{i}|^{2}}{2}\leq\frac{n}{2}\cdot\max_{i\in[k]}|C_{i}|.$
As $e(H)\geq e(G)-\delta_{0}n^{2}\geq pn^{2}/4$, this shows that at least one of the $C_{i}$’s has size at least $pn/2$. Without loss of generality, say $|C_{1}|\geq pn/2$. Then $G[C_{1}]$ has at least $\binom{|C_{1}|}{2}-\delta_{0}n^{2}$ edges, so $G[C_{1}]$ has edge density at least $1-\delta_{0}n^{2}/\binom{|C_{1}|}{2}\geq 1-\delta_{0}n^{2}/(|C_{1}|^{2}/4)\geq 1-\delta_{0}n^{2}/(p^{2}n^{2}/16)=1-\delta$, as desired. ∎
@@
Let $\gamma\in(0,1/30)$, $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph such that $\operatorname{surp}(G)\leq n^{1+\gamma}$. Then $G$ is $\delta$-close to a disjoint union of cliques.
@@
Note that lemma˜5.2 implies $\operatorname{surp}^{*}(G)=O(\operatorname{surp}(G)\log n)\leq n^{1+\gamma^{\prime}}$ for some constant $\gamma^{\prime}\in(\gamma,1/30)$. One of the key differences compared to the proof of Theorem˜1.4 is that we use Lemma 6.6 to have $4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}$ satisfied for every $T\geq n^{1-\frac{1}{24}+\frac{\gamma^{\prime}}{4}}$. Then, setting $q=1/30$, we know that $\gamma^{\prime}<q$ and $1-q>1-\frac{1}{24}+\frac{\gamma^{\prime}}{4}$. So we can apply Lemma˜6.9. Another difference is that we bound $\sum_{0<\lambda_{i}}\lambda_{i}^{2}$ using Lemma˜5.3 (ii), which gives $\sum_{0<\lambda_{i}}\lambda_{i}^{2}\leq O(\sqrt{n}\operatorname{surp}^{*}(G))\leq O(n^{3/2+\gamma^{\prime}})=o(n^{2})$. ∎
@@
Let $\gamma\in(0,1/30)$, $p>0$ and $\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph of edge density $p$ such that $\operatorname{surp}(G)\leq n^{1+\gamma}$. Then $G$ contains a subgraph on at least $pn/2$ vertices of edge density at least $1-\delta$.
@@ 9 Densification — Phase 3
[…]
Theorem 9.1.
For sufficiently large $n$, any $n$-vertex graph $G$ of density at least $1-10^{-6}$ satisfies the following.
  • (a)
    If the smallest eigenvalue $\lambda_{n}$ satisfies $|\lambda_{n}|\leq n^{1/4}/(\log n)^{4}$, then $G$ contains an induced subgraph on $\Omega(n/\log n)$ vertices whose density is at least $1-\widetilde{O}\big(|\lambda_{n}|^{2}/n\big)$.
  • (b)
    If $\operatorname{surp}(G)\leq n^{6/5}/(\log n)^{6}$, then $G$ contains an induced subgraph on $\Omega(n/\log n)$ vertices whose density is at least $1-\widetilde{O}\big(\operatorname{surp}(G)^{2}/n^{3}\big)$.
We note that both $1-\widetilde{O}(|\lambda_{n}|^{2}/n)$ and $1-\widetilde{O}(\operatorname{surp}(G)^{2}/n^{3})$ are tight up to a poly-logarithmic factor. Indeed, with high probability, the Erdős–Rényi graph G(n,1p)G(n,1-p) has smallest eigenvalue $\Theta((pn)^{1/2})$ and surplus $\Theta(n(pn)^{1/2})$ while no induced subgraph has density much larger than 1p1-p. Hence, Theorem˜9.1 shows that under moderate conditions on the smallest eigenvalue or the surplus, any dense graph must contain a large induced subgraph of density very close to $1$.
We then present the intuition behind this result, in particular focusing on (a). To simplify the setup, assume that $G$ is (n1d)(n-1-d)-regular with $d\leq n/10^{6}$ and has smallest eigenvalue $\lambda_{n}$. Then its complement $\overline{G}$ is $d$-regular with second eigenvalue $|\lambda_{n}|-1$. As discussed in Section˜1.4, the work of Balla [10], and Räty, Sudakov, Tomon [13, 74] asserts that the second eigenvalue of a $d$-regular graph is at least
$|\lambda_{n}(G)|-1=\lambda_{2}(\overline{G})=\Omega\Big(\max\big\{d^{1/3},\min\{d^{1/2},n/d\}\big\}\Big)=\begin{cases}\Omega(d^{1/2})&\mbox{ if }d\leq n^{2/3},\\ \Omega(n/d)&\mbox{ if }d\in[n^{2/3},n^{3/4}],\\ \Omega(d^{1/3})&\mbox{ if }d\in[n^{3/4},(1/2-\varepsilon)n].\end{cases}$
Interestingly, this lower bound $f(d)=\Omega(\max\{d^{1/3},\min\{d^{1/2},n/d\}\})$ is not monotone in $d$. However, if
$|\lambda_{n}(G)|\ll n^{1/4}\approx\min_{n^{2/3}\leq d\leq(1/2-\varepsilon)n}f(d),$
then we must have $d\leq n^{2/3}$, showing that $|\lambda_{n}(G)|\geq\Omega(d^{1/2})$, or equivalently, $d\leq O(|\lambda_{n}|^{2})$. Hence, $G$ has density $1-\frac{d}{n}\geq 1-O(\frac{|\lambda_{n}|^{2}}{n})$, as required.
Unfortunately, there are several difficulties to deal with graphs that are not regular, which requires significant new ideas. First, we extend the main results of [74] on eigenvalues and surplus to graphs that are somewhat regular; this is presented in Sections˜9.2 and 9.3. To do this, we employ a novel trick which uses triple Hadamard products. This allows us to complete the proof in Section˜9.4.
9.1 Finding balanced subgraphs
We begin by passing to an induced subgraph of $G$ whose complement is somewhat regular. We say a graph $G$ is $C$-balanced if $\Delta(G)\leq Cd(G)$.
We note that the problem of finding large $C$-balanced (or $C$-almost-regular) induced subgraphs in general graphs was considered by Alon, Krivelevich and Sudakov in [7]. Lemma 9.2 is similar in spirit to their results, but it controls the density of the resulting graph explicitly, which will be useful later.
@@ 9.1 Finding balanced subgraphs
Let $G$ be an $n$-vertex graph of edge density $p\leq 1/5$. Then, $G$ has an induced subgraph $G^{\prime}$ on $\Omega(n/\log n)$ vertices such that the density of $G^{\prime}$ is $p^{\prime}\leq p$, and $G^{\prime}$ is $C$-balanced with $C=4\log_{2}1/p^{\prime}$.
@@
We may assume that $n$ is sufficiently large as otherwise, we can simply take $G^{\prime}$ to be a single vertex. Let $G_{0}=G$ and define the sequence of induced subgraphs $G_{0}\supset G_{1}\supset$... as follows. If the graph $G_{i}$ has $n_{i}$ vertices and density $p_{i}$, and it contains an induced subgraph $G_{i+1}$ on at least $\big(1-\frac{1}{\log_{2}1/p_{i}}\big)n_{i}$ vertices and density $p_{i+1}=p(G_{i+1})<p_{i}/2$, then pick this induced subgraph to be the next element of the sequence (if there are several such subgraphs $G_{i+1}$, choose one arbitrarily). If there is no such subgraph, terminate the procedure.
Suppose that subgraphs $G_{0}\supset G_{1}\supset\dots\supset G_{k}$ have been defined in this way. Then for each $i\leq k$, we have $p_{i}\leq p/2^{i}$, and therefore the process terminates in at most $k\leq 2\log_{2}n$ steps (since we cannot have $p_{k}\leq p/2^{k}\leq 1/n^{2}$ unless $p_{k}=0$, in which case the process terminates). In addition, $\log_{2}1/p_{i}\geq i+\log_{2}1/p$, so $n_{i+1}\geq\big(1-\frac{1}{i+\log_{2}1/p}\big)n_{i}$, which leads to the conclusion
$n_{k}\geq n\prod_{i=0}^{k-1}\left(1-\frac{1}{i+\log_{2}1/p}\right)=\frac{\log_{2}1/p-1}{k-1+\log_{2}1/p}\cdot n\geq\frac{n}{k+1}=\Omega\left(\frac{n}{\log n}\right).$
Finally, to define $G^{\prime}$, remove from $G_{k}$ all vertices of degree at least $(n_{k}-1)p_{k}\log_{2}1/p_{k}$; there are at most $\frac{n_{k}}{\log_{2}1/p_{k}}$ such vertices. Hence, the number of vertices in $G^{\prime}$ is $n^{\prime}\geq\big(1-\frac{1}{\log_{2}1/p_{k}}\big)n_{k}$ and the density of $G^{\prime}$ is $p^{\prime}\leq p_{k}$. Using that $p_{k}\leq p\leq 1/5$, we know that $n^{\prime}\geq(n_{k}+1)/2$. Moreover, since the process terminated at $G_{k}$, no induced subgraph of $G_{k}$ on at least $\big(1-\frac{1}{\log_{2}1/p_{k}}\big)n_{k}$ vertices can have density less than $p_{k}/2$, so $p_{k}/2\leq p^{\prime}\leq p_{k}$. Since the function $x\log_{2}\frac{1}{x}$ is increasing in $(0,1/e)$, we have $\frac{1}{2}p_{k}\log_{2}\frac{2}{p_{k}}\leq p^{\prime}\log_{2}\frac{1}{p^{\prime}}$, thereby $p_{k}\log_{2}\frac{1}{p_{k}}\leq 2p^{\prime}\log_{2}\frac{1}{p^{\prime}}$. Then, the maximum degree of $G^{\prime}$ satisfies
$\Delta(G^{\prime})\leq(n_{k}-1)p_{k}\log_{2}(1/p_{k})\leq 2(n^{\prime}-1)\cdot 2p^{\prime}\log_{2}(1/p^{\prime})=d(G^{\prime})\cdot 4\log_{2}(1/p^{\prime}).$
This completes the proof as $n^{\prime}\geq(n_{k}+1)/2=\Omega(n/\log n)$. ∎
We also use the fact that the density of a $C$-balanced graph is robust under deleting few vertices.
@@
Let $G$ be an $n$-vertex graph with density $p$ that is $C$-balanced for some $C\geq 1$. Then every induced subgraph on at least (11/4C)n(1-1/4C)n vertices has density at least $p/2$.